Stacey brought 45 six cents stamps and eight cents stamps . She paid 3.00 for them. How many of each did she buy? rickeycol@yahoo.com
Answer by mananth(16949)

x 6 cent stamps
y 8 cent stamps

6 x + 8 y = 300 .............1

1 x + 1 y = 45 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 8
-6 x -8 y = -300
8 x + 8 y = 360
Add the two equations
2 x = 60
/ 2
x = 30
plug value of x in (1)
6 x + 8 y = 300
180 + 8 y = 300
8 y = 300 -180
8 y = 120
y = 15
30 6 cent stamps
15 8 cent stamps
m.ananth@hotmail.ca